Q: If ∫x 3 (√ ((1+x 2 )/(1-x 2 )))dx= A √(1-x 4 ) - Bx 2 (√(1-x 4 ) + C sin -1 (x 2 ) + D ,then A+B+C=? A. -1/2 B.-1/4 C.1/4 D.1/2

Q: If ∫x3(√ ((1+x2)/(1-x2)))dx= A √(1-x4) - Bx2(√(1-x4) + C sin-1 (x2) + D ,then A+B+C=?


A. -1/2


B.-1/4


C.1/4


D.1/2

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

∫x3(√ ((1+x2)/(1-x2)))dx= A √(1-x4) - Bx2(√(1-x4) + C sin-1 (x2) + D

differentiate w.r.t x

 and simplify

 x2 +x4 = x2(-2A) + x4(4B) +2C -2B

compair coefficient of same power of x

 -2A =1

  A =-1/2


4B =1

 B=1/4


2C -2B =0

 C=B=1/4


now  A+B+C =0

please check the question again ,answer is not mathcing.

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin